# Urban4D: Semantic-Guided 4D Gaussian Splatting for Urban Scene Reconstruction

### v1

@@ Abstract
Reconstructing dynamic urban scenes presents significant challenges due to their intrinsic geometric structures and spatiotemporal dynamics. Existing methods that attempt to model dynamic urban scenes without leveraging priors on potentially moving regions often produce suboptimal results. Meanwhile, approaches based on manual 3D annotations yield improved reconstruction quality but are impractical due to labor-intensive labeling. In this paper, we revisit the potential of 2D semantic maps for classifying dynamic and static Gaussians and integrating spatial and temporal dimensions for urban scene representation. We introduce Urban4D, a novel framework that employs a semantic-guided decomposition strategy inspired by advances in deep 2D semantic map generation. Our approach distinguishes potentially dynamic objects through reliable semantic Gaussians. To explicitly model dynamic objects, we propose an intuitive and effective 4D Gaussian splatting (4DGS) representation that aggregates temporal information through learnable time embeddings for each Gaussian, predicting their deformations at desired timestamps using a multilayer perceptron (MLP). For more accurate static reconstruction, we also design a k-nearest neighbor (KNN)-based consistency regularization to handle the ground surface due to its low-texture characteristic. Extensive experiments on real-world datasets demonstrate that Urban4D not only achieves comparable or better quality than previous state-of-the-art methods but also effectively captures dynamic objects while maintaining high visual fidelity for static elements.
@@
Figure 2: Different from previous 3DGS-based methods which leverage pre-trained segmentation models to predict 2D semantic maps for downstream tasks, we revisit the potential of these estimated semantic maps to handle potentially dynamic objects for reconstruction. From left to right: input image, semantic map, our reconstruction, and reconstruction without potentially dynamic objects. Our method effectively preserves the geometry and appearance of both static structures and potentially dynamic objects.
@@ 1 Introduction
Recent advancements in urban scene reconstruction have witnessed remarkable progress, particularly with the introduction of 3D Gaussian Splatting (3DGS)  [13], which has enhanced real-time rendering while maintaining high reconstruction quality for static scenes. However, directly extending these capabilities to model dynamic and large-scale urban scenes presents significant challenges due to spatiotemporal dynamics and intricate geometric structures.
Urban scenes can generally be categorized into two types: dynamic objects, such as vehicles and pedestrians, which exhibit complex motion patterns, and static elements, which maintain consistent spatial relationships. Specifically, low-texture regions (e.g., ground surfaces) pose challenges for accurate static reconstruction due to their lack of salient patterns. These unique characteristics make reconstructing dynamic objects in urban scenes particularly difficult. The coexistence of static and dynamic elements necessitates a model that can distinguish between and process them independently to ensure accurate reconstruction.
Previous dynamic NeRF-based methods, such as  [38, 32], attempt to address these challenges by introducing time-conditioned neural networks to model motion, but they suffer from slow rendering speeds and limited capacity to handle complex and large-scale urban scenes. Some methods explore suitable representations, such as Periodic Vibration Gaussian (PVG)  [7], which proposes a unified framework by incorporating periodic vibration-based temporal dynamics in urban scenes. However, such a unified representation may degrade performance when treating static elements with temporal processes. Other 3DGS-based methods resort to manually labeled 3D bounding boxes to process dynamic objects  [37, 41], which is labor-intensive and impractical for large-scale or continuously changing environments.
Furthermore, we have witnessed that some of these methods adopt pre-trained segmentation models, such as SegFormer [35], for semantic novel view rendering. Semantic segmentation based on deep learning provides more reliable classifications of scene components in 2D images. Leveraging these results, we may effectively classify dynamic and static Gaussians without relying on manually labeled 3D annotations. The observations in Figure 2 lead us to an intriguing question: Can we revisit the potential of 2D semantic maps for classifying potentially dynamic objects and leverage spatiotemporal information to adequately represent urban scenes? The answer to this question uncovers an unexplored area in reconstructing dynamic objects based on 4D Gaussian splatting (4DGS), enabling a more holistic and precise representation.
Inspired by this perspective, we propose Urban4D, a novel framework to overcome the limitations of traditional 3DGS-based methods in handling temporal dynamics and reconstructing urban scenes. Specifically, we design a semantic-guided decomposition strategy, which builds on recent progress in deep 2D semantic map generation to distinguish potentially dynamic objects from static elements by assigning reliable semantic classes to Gaussians. To explicitly model dynamic objects, we propose an intuitive and effective 4DGS representation of dynamic objects by incorporating a learnable time embedding with each Gaussian and predicting Gaussian deformations at desired timestamps through a multilayer perception (MLP). Moreover, we employ a k-nearest neighbor (KNN)-based consistency regularization mechanism to enforce consistency across neighboring Gaussians to alleviate the inaccurate static reconstruction caused by the low-texture characteristic of the ground surface.
Our extensive experiments on multiple real-world datasets demonstrate that Urban4D achieves comparable or better quality than previous state-of-the-art methods. The proposed framework effectively captures dynamic objects while maintaining high visual fidelity for static elements. By addressing the limitations of existing methods and introducing innovative techniques for semantic-guided decomposition and temporal modeling, Urban4D appears to be a compelling solution for representing the complexity and detail of dynamic urban environments. Our contributions can be summarized as follows:
@@ 1. $\bullet$
    We propose Urban4D, a novel framework that extends the capabilities of 3DGS to effectively handle dynamic urban scenes by leveraging semantic and spatiotemporal information.
@@ 2. $\bullet$
    An intuitive semantic-guided decomposition strategy based on 2D semantic maps, which distinguishes between dynamic and static elements through reliable semantic Gaussians.
@@ 3. $\bullet$
    An intuitive and efficient 4D Gaussian splatting representation that incorporates a learnable time embedding for each Gaussian and predicts deformations over time with an MLP, effectively modeling dynamic objects in urban scenes.
  4. $\bullet$
    A KNN-based consistency regularization mechanism to handle the low-texture ground surface in urban scenes, improving the accuracy of static reconstruction.
@@ 2 Related Work
2.1 Neural Scene Representations
Neural representations have revolutionized novel view synthesis, with NeRF [20] leading significant advances in this field. NeRF utilizes multi-layer perceptrons (MLPs) and differentiable volume rendering to reconstruct 3D scenes from 2D images and camera poses. While demonstrating impressive results for bounded scenes, its application to large-scale unbounded scenes remains challenging due to computational constraints and the requirement for consistent camera-object distances.
@@ 2.2 Dynamic Scene Modeling
To address these challenges, recent works have proposed various scene decomposition strategies. Scene graph representations [39, 32, 37, 41, 27] enable explicit object-level modeling and control. However, most current approaches either treat all dynamic elements uniformly [7] or rely heavily on manual annotations [41, 37].
@@ 3 Methodology
Method overview. Our proposed method aims to reconstruct dynamic urban scenes by leveraging semantic information to effectively distinguish between static and dynamic elements.
Given a sequence of images $\{I_{t}\}_{t=1}^{T}$ and corresponding LiDAR point clouds $\{P_{t}\}_{t=1}^{T}$ captured by a moving vehicle, we aim to reconstruct dynamic urban scenes. For each frame, semantic maps $\{S_{t}\}_{t=1}^{T}$ are predicted using an off-the-shelf pre-trained segmentation model. Based on our aforementioned insight, we develop a novel framework to intuitively represent dynamic elements in 4D and enhance the reconstruction of static elements.
Our method consists of three main elements (Figure 4): (1) semantic-guided decomposition that separates the scene into static and potentially dynamic Gaussians based on semantic information (Sec. 3.2), (2) 4D Gaussians Splatting representation for dynamic scenes(Sec. 3.3), and (3) a regularization method that enforces consistency in low-texture regions, particularly ground surfaces, through KNN-based optimization (Sec. 3.4).
@@ 3.1 Preliminaries
Deformable Gaussian Splatting represents a dynamic scene using a set of deformable 3D Gaussians $\{\mathcal{G}_{i}\}_{i=1}^{N}$. Each Gaussian $\mathcal{G}_{i}$ is characterized by its canonical position $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, rotation $\mathbf{r}_{i}\in\mathbb{R}^{4}$, and scale $\mathbf{s}_{i}\in\mathbb{R}^{3}$. At time $t$, both position and shape undergo deformation:
$\boldsymbol{\mu}_{i}(t)$ $=\boldsymbol{\mu}_{i}+\mathcal{F}_{pos}(\boldsymbol{\mu}_{i},t),$ (1)
$\mathbf{r}_{i}(t)$ $=\mathbf{r}_{i}+\mathcal{F}_{rot}(\boldsymbol{\mu}_{i},t),$
$\mathbf{s}_{i}(t)$ $=\mathbf{s}_{i}+\mathcal{F}_{scale}(\boldsymbol{\mu}_{i},t),$
where $\mathcal{F}_{pos}$, $\mathcal{F}_{rot}$, and $\mathcal{F}_{scale}$ are deformation MLPs predicting position offset, rotation change, and scale adjustment respectively based on the Gaussian’s position and time. The covariance matrix at time $t$ is then computed as:
$\boldsymbol{\Sigma}_{i}(t)=\mathbf{r}_{i}(t)\mathbf{s}_{i}(t)\mathbf{s}_{i}(t)^{T}\mathbf{r}_{i}(t)^{T}.$ (2)
Each Gaussian also maintains its opacity $\alpha_{i}\in[0,1]$ and color represented as spherical harmonics coefficients $\mathbf{c}_{i}$ for appearance modeling. The rendering process follows standard Gaussian splatting with the deformed parameters.
@@ 3.2 Semantic-guided Decomposition
Effectively separating dynamic and static elements is essential for handling urban scenes. Leveraging 2D semantic maps from a pre-trained segmentation model, we introduce a systematic approach to identify and separate potentially dynamic elements, allowing for specialized treatment of movable objects and static elements. Specifically, as shown in Figure 3, for every timestamp, each Gaussian $\mathcal{G}_{i}$ of the current frame is assigned a semantic label $s_{i}\in\{1,...,K\}$, obtained through rendering and supervision of semantic maps, where $K$ represents the total number of semantic classes.
Our method focuses on potentially dynamic objects in urban scenes, typically comprising well-defined semantic categories. We identify dynamic classes $\mathcal{C}_{d}$, including vehicles, pedestrians, and cyclists, which frequently exhibit motion and require specialized handling. The remaining classes form the static set $\mathcal{C}_{s}$, including buildings, vegetation, and road surfaces. This semantic understanding enables the decomposition of the scene into two disjoint sets:
@@
where $\mathcal{G}_{i}^{d}$ represents dynamic Gaussians (i.e., Gaussians of potentially dynamic objects) requiring 4D modeling, and $\mathcal{G}_{i}^{s}$ represents static Gaussians. This semantic-guided decomposition offers several key advantages: (1) it enables specialized motion modeling for dynamic objects while preserving static scene consistency, (2) it eliminates the need for labor-intensive manual annotations, and (3) it ensures consistent motion treatment for all dynamic objects, regardless of their movement state.
The decomposition lays the foundation for our two-stream optimization strategy. Dynamic Gaussians in $\mathcal{G}_{i}^{d}$ undergo dedicated motion refinement (Sec. 3.3) to accurately capture movement, while static Gaussians receive geometric regularization (Sec. 3.4) to enhance scene stability. Additionally, we employ an optimizable environment texture map for sky representation, which is rendered separately and composited with the Gaussian-rendered image through alpha blending, as described in [7]. This separation allows targeted optimization without interference between dynamic and static components. Optimizing stationary dynamic Gaussians (e.g., parked vehicles) within a distinct subset is still more efficient than methods like DeformGS [40], which jointly model entire scene dynamics. This targeted approach results in a more accurate reconstruction of both dynamic and static elements.
Figure 4: Overview of Urban4D framework. Given input images with semantic information, Gaussians are classified into static and potentially dynamic elements through semantic-guided decomposition. Dynamic objects are represented in 4D Gaussian Splatting that captures motion patterns by incorporating a learnable time embedding, with deformations predicted at desired timestamps using an MLP. For static Gaussians, particularly low-texture areas like ground surfaces, consistency is enforced using a KNN-based regularization mechanism.
@@
To represent our dynamic Gaussians $\{\mathcal{G}_{i}^{d}\}$ in 4D, our method extends the deformation mechanism of DeformGS [40] by introducing a learnable time embedding for each Gaussian. While DeformGS [40] directly maps spatial positions and time to deformations, our approach leverages temporal context through Gaussian-specific embeddings.
@@
3.4 Ground Surface Consistency Regularization
In urban driving scenes, ground surfaces constitute a significant portion of the environment and typically exhibit low-texture characteristics. While ground-level Gaussians should theoretically share similar properties due to their homogeneous nature, enforcing strict uniformity across all ground Gaussians would be oversimplified and impractical, as real-world surfaces often contain variations and irregularities. The scale parameter of a Gaussian, derived from its covariance matrix, inherently encodes local geometric information analogous to surface normals [12, 8]. A well-behaved ground surface should exhibit smooth transitions in its local geometry, making scale a particularly suitable target for regularization. This motivates us to regularize the scale parameters rather than other Gaussian properties.
$\mathcal{N}_{i}=\text{KNN}(\mathcal{G}_{i},\mathcal{G}_{g},N),$ (7)

### v2

# UrbanGS: Semantic-Guided Gaussian Splatting for Urban Scene Reconstruction

@@ Abstract
Reconstructing urban scenes is challenging due to their complex geometries and the presence of potentially dynamic objects. 3D Gaussian Splatting (3DGS)-based methods have shown strong performance, but existing approaches often incorporate manual 3D annotations to improve dynamic object modeling, which is impractical due to high labeling costs. Some methods leverage 4D Gaussian Splatting (4DGS) to represent the entire scene, but they treat static and dynamic objects uniformly, leading to unnecessary updates for static elements and ultimately degrading reconstruction quality. To address these issues, we propose UrbanGS, which leverages 2D semantic maps and an existing dynamic Gaussian approach to distinguish static objects from the scene, enabling separate processing of definite static and potentially dynamic elements. Specifically, for definite static regions, we enforce global consistency to prevent unintended changes in dynamic Gaussian and introduce a K-nearest neighbor (KNN)-based regularization to improve local coherence on low-textured ground surfaces. Notably, for potentially dynamic objects, we aggregate temporal information using learnable time embeddings, allowing each Gaussian to model deformations over time. Extensive experiments on real-world datasets demonstrate that our approach outperforms state-of-the-art methods in reconstruction quality and efficiency, accurately preserving static content while capturing dynamic elements.
@@ 1 Introduction
Urban scenes are characterized by two primary categories of objects: major static elements, including buildings and road infrastructure, which remain spatially consistent over time, and some potentially dynamic elements, such as pedestrians and vehicles, which can remain static or exhibit diverse and often unpredictable motion patterns. Accurate reconstruction of urban scenes thus remains challenging, mainly due to the coexistence of these static and dynamic elements and complexities arising from low-textured regions.
Recent advancements in 3D Gaussian Splatting (3DGS)  [43, 48, 8, 10], which have attempted to incorporate manually labeled 3D bounding boxes to process dynamic objects separately. However, such manual annotations are labor-intensive, impractical for large-scale settings, and unsuitable for continuously evolving environments where dynamic objects frequently change positions. Alternative approaches leverage 4DGS-based representations, such as Periodic Vibration Gaussian (PVG) [7], which introduces periodic temporal modeling to represent motion variations in urban scenes. However, these methods lack explicit differentiation between static and dynamic elements, leading to unnecessary updates for stationary objects, which in turn degrades reconstruction quality.
We observe that Gaussians supervised by semantic maps inherently acquire semantic information, which can be leveraged for identifying static objects. Furthermore, deep-learning-based 2D semantic segmentation models provide robust classification capabilities, allowing Gaussians to be categorized into determined static and potentially dynamic elements without relying on explicit 3D annotations.
Inspired by this perspective, we introduce UrbanGS, a semantic-guided Gaussian Splatting framework designed to effectively handle definite static elements in urban scene reconstruction while adapting to potentially dynamic components. Specifically, for definite static Gaussians, we introduce a global consistency constraint to ensure that they remain unchanged over time. Additionally, we employ a K-nearest neighbor (KNN)-based consistency regularization to improve local coherence, particularly for low-textured surfaces such as roads, which pose significant challenges in urban scene reconstruction. For potentially dynamic Gaussians, we propose an efficient 4DGS representation that incorporates learnable time embeddings for each Gaussian. This design enables the model to predict object deformations at arbitrary timestamps using a lightweight multilayer perceptron (MLP), effectively capturing urban dynamics while preserving rendering efficiency.
Our extensive experiments on real-world urban datasets demonstrate that UrbanGS achieves state-of-the-art reconstruction quality in both static and potentially dynamic objects. The key contributions of this work can be summarized as follows:
@@ 1. $\bullet$
    We introduce UrbanGS, a novel semantic-driven framework that leverage 2D semantic segmentation to separate static Gaussians from potentially dynamic Gaussians without requiring manual 3D annotations.
@@ 2. $\bullet$
    We propose a global consistency constraint to enforce temporal stability in static Gaussians, preventing unnecessary updates and significantly improving reconstruction quality. Additionally, to address the challenge of low-textured regions, we introduce a KNN-based consistency regularization, ensuring a more stable and accurate reconstruction of surfaces such as roads and sidewalks.
@@ 3. $\bullet$
    We develop a learnable time embedding mechanism for potentially dynamic Gaussians, enabling the model to predict object deformations at arbitrary timestamps using a lightweight MLP-based deformation model, efficiently handling motion in urban scenes.
@@ 2 Related Work
Neural Scene representations have revolutionized novel view synthesis, with NeRF [24] leading significant advances in this field. NeRF utilizes multi-layer perceptrons (MLPs) and differentiable volume rendering to reconstruct 3D scenes from 2D images and camera poses. While demonstrating impressive results for bounded scenes, its application to large-scale unbounded scenes remains challenging due to computational constraints and the requirement for consistent camera-object distances.
@@
To address these challenges, recent works have proposed various scene decomposition strategies. Scene graph representations [45, 38, 43, 48, 32, 8, 10] enable explicit modeling and control at the object level. However, most current approaches either treat all dynamic elements uniformly [7, 46, 15] or rely heavily on manual annotations [48, 43, 8, 10].
Gaussian reconstruction with semantic features Recent advances integrate 3DGS with semantic features. Feature 3DGS [47] extends 3D Gaussian Splatting by introducing high-dimensional feature fields. Similarly, Semantic Gaussians [13] tackles open-vocabulary 3D scene understanding by mapping diverse 2D semantic features into 3D Gaussian. These works focus on static scenes, while our work complements these efforts by focusing on urban scenes, leveraging semantic decomposition for static/dynamic separation.
@@ 3 Methodology
Figure 3: Overview of UrbanGS framework. Given input images with semantic information during training, Gaussians are classified into definite static and potentially dynamic elements through semantic-guided decomposition. For definitively static Gaussians, we introduce a static invariance constraint to preserve their temporal invariance and prevent unintended transformations. To address challenges in low-texture regions (e.g., ground surfaces), a KNN-based regularization mechanism is employed to enforce structural coherence. Potentially dynamic objects are represented in 4D Gaussian Splatting that captures motion patterns by incorporating a learnable time embedding, with deformations predicted at desired timestamps using an MLP.
Method overview. Our proposed method aims to reconstruct dynamic urban scenes by leveraging semantic information to effectively distinguish between determined static and potentially dynamic elements.
Given a sequence of images $\{I_{t}\}_{t=1}^{T}$ and the corresponding LiDAR point clouds $\{P_{t}\}_{t=1}^{T}$ captured by a moving vehicle, our aim is to reconstruct urban scenes. For each frame, semantic maps $\{S_{t}\}_{t=1}^{T}$ are predicted using an off-the-shelf pre-trained segmentation model. Building upon a uniformly dynamic Gaussian approach, which inherently applies time-dependent transformations to the global Gaussians, such as PVG [7], we propose a novel framework that introduces semantic-aware improvement. Specifically, during training, we leverage the semantic attributes of each Gaussian to enforce constraints and adaptively adjust the properties of each Gaussian point. This allows us to effectively keep static elements unchanged across time, enforce consistency in low-texture regions, and intuitively capture potentially dynamic objects through a 4D representation. Our approach enhances the robustness and accuracy of urban scene reconstruction by combining the strengths of semantic guidance and dynamic Gaussian modeling. More details of pseudo algorithm for training are included in the supplementary material.
Our method consists of three main elements (Figure 3): (1) semantic-guided decomposition that separates the scene into static and potentially dynamic Gaussians based on semantic information (Sec. 3.2), (2) a temporal-invariance regularization for all static points, ensuring they remain unchanged over time, and apply a KNN-based consistency constraint to low-texture regions for enhanced reconstruction fidelity (Sec. 3.3), and (3) a 4d Gaussians Splatting representation for potentially dynamic objects (Sec. 3.4) .
@@ 3.1 Preliminaries
3D Gaussian Splatting represents a scene as a set of 3D Gaussians $\{\mathcal{G}_{i}\}_{i=1}^{N}$, where each Gaussian $\mathcal{G}_{i}$ is parameterized by its mean position $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, covariance matrix $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$, and appearance features including opacity $\alpha_{i}\in\mathbb{R}$ and spherical harmonics coefficients $\mathbf{f}_{i}\in\mathbb{R}^{48}$ for RGB color representation.
For each pixel in the target view, the rendering process involves projecting the 3D Gaussians onto the 2D image plane. The projection of a 3D Gaussian results in a 2D Gaussian with parameters:
$\boldsymbol{\mu}_{2D}=\Pi(\boldsymbol{\mu}_{i}),$ (1)
$\boldsymbol{\Sigma}_{2D}=J\boldsymbol{\Sigma}_{i}J^{T},$ (2)
where $\Pi(\cdot)$ is the perspective projection function and $J$ is the Jacobian of the projection.
The final color $C(x,y)$ at pixel $(x,y)$ is computed through alpha compositing:
$C(x,y)=\sum_{i=1}^{N}T_{i}\alpha_{i}c_{i},$ (3)
where $T_{i}$ represents the accumulated transmittance, $\alpha_{i}$ is the opacity, and $c_{i}$ is the RGB color from spherical harmonics coefficients. The scene is optimized by minimizing the difference between rendered and ground-truth images.
In addition to RGB supervision, each Gaussian can be associated with semantic labels $s_{i}\in\{1,...,K\}$, where $K$ is the number of semantic classes. The semantic prediction at pixel $(x,y)$ can be computed similarly through alpha compositing:
$S(x,y)=\sum_{i=1}^{N}T_{i}\alpha_{i}s_{i},$ (4)
@@ 3.2 Semantic-guided Decomposition
Leveraging 2D semantic maps from a pre-trained segmentation model, we introduce a systematic approach to identify and separate static elements, allowing for specialized treatment of static elements and remaining potentially movable objects. Specifically, as shown in Figure 2, during dynamic Gaussian training, each Gaussian point $\mathcal{G}_{i}$ is assigned a semantic label $s_{i}\in\{1,...,K\}$, obtained through rendering and supervision of semantic maps, where $K$ represents the total number of semantic classes. Our method focuses on determined static objects in urban scenes, typically comprising well-defined semantic categories such as building, roads and trees. We have designed a strategy specifically for those static objects to ensure their invariance over time. We also identify potentially dynamic classes $\mathcal{C}_{d}$, including vehicles, pedestrians, and cyclists, which frequently exhibit motion and require specialized handling. This semantic understanding enables the decomposition of the scene into two disjoint sets:
@@
where $\mathcal{G}_{i}^{s}$ denotes the static Gaussians that require static regularization, while $\mathcal{G}_{i}^{d}$ refers to dynamic Gaussians (i.e., Gaussians associated with potentially dynamic objects) that necessitate 4D modeling. This semantic-guided decomposition offers several key advantages: (1) it ensures static elements remain unchanged over time while confining temporal modeling to potentially dynamic objects, (2) it enhances the reconstruction quality specifically for road, and (3) it eliminates the need for labor-intensive manual annotations.
The decomposition lays the foundation for our two-stream optimization strategy. the static Gaussians receive geometric regularization (Sec. 3.3) to enhance scene stability, while the potentially dynamic Gaussians in $\mathcal{G}_{i}^{d}$ undergo a dedicated motion refinement (Sec. 3.4) to accurately capture movement. Additionally, we employ an optimizable environment texture map for sky representation, which is rendered separately and combined with the Gaussian-based image by alpha blending, as described in [7].
3.3 Static Regularization
Static invariance. Prior approaches (e.g., PVG) address dynamic scene reconstruction by applying timestamp-dependent transformations to each Gaussian’s 3D position $\mu$ and opacity $\alpha$. These transformations effectively capture dynamic motions but inevitably alter truly static parts of the scene. To mitigate this issue, we introduce a static consistency loss to keep these static Gaussians invariant:
$\mathcal{L}_{\text{static}}=\sum_{i\in\mathcal{G}_{i}^{s}}w_{i}\Bigl{(}\|\mu_{i}-\mu_{i}^{\prime}\|^{2}+\|\alpha_{i}-\alpha_{i}^{\prime}\|^{2}\Bigr{)},$
where $\mu_{i}$ and $\alpha_{i}$ denote the untransformed parts of static Gaussians, and ${\mu_{i}^{\prime}}$ and $\alpha_{i}^{\prime}$ are their transformed counterparts. $w_{i}$ is a semantic weight (derived from a softmax ratio) indicating the likelihood that Gaussian $i$ is truly static. This weighting mechanism allows fully static points ($w_{i}\approx 1$) to remain nearly unchanged, while points that are partly dynamic or uncertain ($10<w_{i}<$ 1) retain the freedom to move if necessary.
Ground surface consistency regularization. In urban driving scenes, ground surfaces constitute a significant portion of the environment and typically exhibit low-texture characteristics. While ground-level Gaussians should theoretically share similar properties due to their homogeneous nature, enforcing strict uniformity across all ground Gaussians would be oversimplified and impractical, as real-world surfaces often contain variations and irregularities. The scale parameter of a Gaussian, derived from its covariance matrix, inherently encodes local geometric information analogous to surface normals [16, 9]. A well-behaved ground surface should exhibit smooth transitions in its local geometry, making scale a particularly suitable target for regularization. This motivates us to regularize the scale parameters rather than other Gaussian properties.
$\mathcal{N}_{i}=\text{KNN}(\mathcal{G}_{i},\mathcal{G}_{g},N),$ (6)
@@ 3.3 Static Regularization
While the original dynamic Gaussian approach demonstrates the capability to model dynamic objects, we aim to refine it further to better handle the remaining potentially dynamic Gaussians. To refine these potentially dynamic Gaussians ${\mathcal{G}_{i}^{d}}$ in 4D, our method extends the deformation mechanism of DeformGS [46] by introducing a learnable time embedding for each Gaussian. Unlike DeformGS [46], which directly maps spatial positions and time to deformations, our approach leverages temporal context through Gaussian-specific embeddings. This refinement enables a more adaptive representation of dynamic elements.